# arXiv:2603.00321v1[hep-th] 27 Feb 2026

### v1

@@ Abstract
We establish a universal relation between the coefficient $C_{T}$ of the energy–momentum tensor two‑point function and the coefficient $c$ multiplying the term quadratic in the Weyl tensor in the Weyl anomaly of a generic even‑dimensional conformal field theory. Our first derivation combines long‑known holographic results for $C_{T}$ and for the Weyl anomaly in Einstein bulk gravity with a recently obtained Chern–Gauss–Bonnet formula for compact Einstein manifolds. This theorem isolates the Weyl‑squared contribution in the relation between the Euler density and the $Q$-curvature, allowing us to identify the relevant quadratic term unambiguously. We then provide a genuine CFT derivation based on the renormalization‑group running of the TT correlator with respect to the arbitrary but necessary mass scale $\mu$. Several known examples are revisited to illustrate and validate the general result.
@@ I Introduction
In even spacetime dimensions, CFTs exhibit a trace (or Weyl or conformal) anomaly when coupled to a background metric[2]. Although the classical theory is Weyl invariant, the quantum effective action acquires a non‑vanishing trace proportional to local curvature invariants. The anomaly takes the schematic form[3]
$(4\pi)^{d/2}\langle T^{\mu}_{\mu}\rangle=-a\,E_{d}+\sum_{i}c_{i}I_{i}+ttd,$ (2)
where the Euler density $E_{d}$ carries the type-A central charge $a$, the pointwise Weyl invariants $I_{\mathnormal{i}}$ carry the type-B central charges $c_{\mathnormal{i}}$, and ttdttd stands for trivial anomalies or trivial total derivatives whose conformal primitives are local curvature invariants. The central charges, traditionally denoted $a$ and $c$ in four dimensions and by suitable generalizations in higher even dimensions, are intrinsic CFT data.
@@
The explicit mapping between $C_{T}$ and the anomaly coefficients varies with the dimension $d$ of the CFT:
  • d=2 There is only one central charge, $c$, and it is defined by
    $4\pi\langle T^{\mu}{}_{\mu}\rangle=\frac{c}{6}R.$
@@
    $\langle T^{\mu}{}_{\mu}\rangle=-a\,E_{6}+c_{1}I_{1}+c_{2}I_{2}+c_{3}I_{3}+\text{ttd},$
    where $I_{1}$ and $I_{2}$ are cubic in the Weyl tensor, while $I_{3}$ contains terms of the schematic form $W\nabla^{2}W+lot+ttd$. The lower-order terms in derivatives (lot\ lot) clearly do not contribute to the quadratic expansion around flat space and do not bear any connection to $C_{T}$. Therefore, it is the central charge $c_{3}$ that relates to $C_{T}$ [1, 5]
@@
    In holographic Einstein gravity, where the bulk action is just the Hilbert-Einstein action plus a negative cosmological constant, the coefficients of the holographic Weyl anomaly [6, 7] are all determined by a single bulk coupling (that of the critical $Q$-curvature[8]). That enforces linear relations among the various $c_{i}$ and $a$, but structurally, $C_{T}$ still tracks the unique quadratic-in-$W$ direction in the space of invariants, just as in 6d it tracks $c_{3}$. In $8d$, Chen and Lü[9] have recently found a quadratic in Weyl invariant $I^{(8)}_{10}$ whose coefficient $c_{10}$ becomes equal to $a$ in holographic Einstein gravity. Comparing with the long-known result of Liu and Tseytlin for $C_{T}$[10], they obtain (in our conventions)
@@
    where we have chosen a particular pointwise Weyl invariant $W^{(d)}_{2,1}$, recently introduced by Case et at. in [11] that descends from the ambient $\left(\widetilde{\nabla}^{2}\right)^{d/2-2}\widetilde{W}^{2}$ in the Fefferman-Graham program. Restricted to Einstein manifolds, it enters an explicit expression for the Chern-Gauss-Bonnet formula that allows for the interchange between the Euler density and the $Q$-curvature. With this insight, we tune the coefficient $c$ in the anomaly to enforce the equality $c=a$ when the anomaly becomes just the $Q$-curvature.
    Alternatively, we are also free to choose instead a different Weyl invariant quadratic in the Weyl tensor that naturally extends the $W^{2}$ and the $I_{3}$ from 4d and 6d, respectively (see e.g. [12]). The numerical coefficient is obtained by the expansion of the $Q$-curvature to terms quadratic in curvature (see subsection C below) to enforce the equality $c=a$ for the $Q$-curvature,
@@
In what follows we will explain how to obtain the following universal relation between $C_{T}$ and $c$, that accommodates the low dimensions and extrapolates to any even higher dimension [^1] Dimension two is special, our general result can be modified to include this particular case as well by absorbing a factor six in the definition of the 2d $c$, setting instead $4\pi\langle T^{\mu}{}_{\mu}\rangle=c\,R$
$C_{T}=\dfrac{d}{d-1}\cdot\dfrac{(d+1)!}{(d/2-1)!}\cdot c\qquad\text{ for even }d>2.$ (5)
The present note aims to derive the foreseen universal relation between the coefficient $C_{T}$ of the energy-momentum tensor two-point function and the coefficient $c$ of the term quadratic in the Weyl tensor of the associated Weyl anomaly for a generic even-dimensional CFT.
Understanding these relations is essential for characterizing the space of consistent CFTs in higher dimensions and for connecting flat‑space observables with geometric responses. The present work develops this connection in full generality for even‐dimensional CFTs. We follow a holographic approach and combine the long-known holographic results for $C_{T}$ [10] and Weyl anomaly [6],[7],[14] for bulk Einstein gravity, together with a recently derived Chern-Gauss-Bonnet theorem on a compact Einstein manifold that allows one to pinpoint the term quadratic in the Weyl tensor in the relation between Euler density and $Q$-curvature.
We also confirm our findings through a purely CFT computation that exploits the fact that the trace anomaly determines the scaling behavior of the CFT partition function or the effective action.
We start in Section II.1, by reviewing the holographic computation of $C_{T}$ and relating it to the bulk Newton’s constant, or, equivalently, the Planck length. In II.2, we discuss the holographic Weyl anomaly for Einstein gravity with a negative cosmological constant and its relation with the $Q$-curvature. In II.3, we present the Chern-Gauss-Bonnet theorem on compact even-dimensional Einstein manifolds, as recently obtained by Case et al. [11], which relates the Euler density to the $Q$-curvature and pointwise Weyl invariants. In particular, we keep track of the coefficient $c$ of the quadratic term in the Weyl tensor. In II.4, we combine the previous ingredients to obtain the universal relation between $C_{T}$ and $c$. In Section III, we present an alternative derivation within the CFT. Finally, in Section IV, we compare the universal relation with several known examples, illustrate them, and conclude in Section V.
@@ II.3 Weyl-squared content of the $Q$-curvature
$\mbox{Pf}=(2\lambda)^{d/2}\cdot(d-1)!!\,+\,\dfrac{(-2)^{2-d/2}}{(d/2-1)!}\cdot\frac{1}{8}\,W^{(d)}_{2,1}+\text{lot}+\text{ttd}.$ (12)
@@
$E_{d}=\dfrac{(-)^{d/2}}{2^{d/2}}\,\epsilon_{d}\,\epsilon_{d}\,R...R\text{ and }\mbox{Pf}=\dfrac{(d-1)!!}{2^{d/2}\,d!}\,\epsilon_{d}\,\epsilon_{d}\,R...R\,=\,\dfrac{(-)^{d/2}\,(d-1)!!}{d!}\cdot E_{d}.$ (13)
@@
We are now in a position to write down the holographic value for $c$ from Eqn.(11)
@@
$Q_{d}=-\,\dfrac{d-2}{8(d-3)}\,W(-\nabla^{2})^{d/2-2}W+\text{lot}+\text{ttd}.$ (17)
@@
III Universal relation between $C_{T}$ and $c$ : CFT derivation
The CFT derivation exploits the fact that the trace anomaly controls the scale variation of the partition function $Z_{CFT}$ or effective action $W_{CFT}$ (see e.g. [1])
@@ III Universal relation between $C_{T}$ and $c$ : CFT derivation
On the RHS, the kernel of the second-order variation of the trace anomaly is given. It is insufficient to track the term quadratic in the Weyl tensor. Furthermore, the Euler term does not contribute.
One can directly expand the term quadratic in Weyl up to second order in the metric perturbation around flat space, or trade it for the Euler term and $Q$-curvature modulo total derivatives. The $Q$-curvature is pure Ricci and can be easily varied in 4D and 6D, where explicit expressions for it are at hand. Cubic and higher-order terms in the Weyl tensor can safely be discarded when expanding around flat space in the computation of the two-point function, since their contribution to the effective action is cubic or higher in the metric perturbation.
@@
that must be cancelled by the one-loop counterterm, which in turn, is given by the trace anomaly (see e.g. [22], eqns. 21-26 therein for 2D and 4D  [^5] We are grateful to M.J. Duff for information on this point. This same pole matching led to the discovery of the GJMS operators in the two-point function for scalar operators in [50]. See also  [51] for a thorough analysis of scalar operators..
@@ III.0.4 Alternative derivation in 6D
After some algebraic manipulations while expanding the Ricci tensor and scalar, as expected, one obtains the very same relation as before  [^6] This 6D relation was announced by Sen and Sinha [24], but it contained a wrong numerical coefficient in the holographic computation of the central charge $c_{3}$ that spoiled the relation with $C_{T}$ in the journal version. The arXiv version 4, in turn, contains the correct result that agrees with the holographic expectation; see eqns. (3.22) and (D.29) therein. We are grateful to Felipe Diaz for clarification of this issue..
@@ III.0.6 General even $D$: $Q$-curvature and GJMS-like operators
But there is a remarkable observation following this line of reasoning: the second metric variation of the $Q$-curvature, which can be thought of as a particular conformal gravity Lagrangian admitting Einstein solutions, is just the kinetic term of the corresponding Weyl graviton when restricted to the transverse-traceless component of the metric fluctuation. This is nothing but the GJMS-like operator on symmetric transverse-traceless two-tensors that happens to factorize into shifted Lichnerowicz Laplacians on Einstein manifolds, as shown some time ago by Matsumoto [30]. In particular, in flat space, it reduces to a power of the Laplacian. It is just a matter of keeping track of the overall coefficient to be able to obtain the numerical factor that fixes the RHS of eqn.(32) [^8] We are grateful to A.A.Tseytlin for discussion of this point.
@@ IV Examples
1 $\partial^{2}$ conformal scalar 8/7 6
2 $\partial^{4}$ conformal scalar -32/7 -24
3 $\partial^{6}$ conformal scalar 24 126
4 $\partial^{8}$ conformal scalar -256 -1344
@@
$C^{E}_{T}=4\pi^{d/2}\cdot\frac{\Gamma(d+2)}{(d-1)\,\Gamma^{3}(d/2)}\cdot\frac{L^{d-1}}{l_{\text{P}}^{d-1}}\qquad vs.\qquad c^{E}=\dfrac{4\pi^{d/2}}{d\,\Gamma^{2}(d/2)}\cdot\dfrac{L^{d-1}}{l_{P}^{d-1}}~.$ (42)
@@
Let us follow a shortcut route [39, 40, 41] to get this algebraic equation  [^9] Of course, this is equivalent to extremizing the gravitational action with respect to the putative AdS radius as in [33].. We exploit the scale invariance of the action to write
@@
$S_{\text{G}}[\mu^{2}\hat{g}_{AdS}]=\mu^{d+1}\int_{\mathcal{M}}d^{d+1}x\sqrt{\hat{g}_{AdS}}\;{\cal L}(\frac{f_{\infty}}{\mu^{2}}),$ (46)
@@
We can now apply our recipe to determine the holographic central charge $a$ by evaluating it on a particular Poincaré-Einstein bulk metric with an Einstein metric at its conformal infinity. From the pure-Ricci part, we get the action at $AdS$ precisely, and from the deviations, involving the Weyl tensor, it is enough to keep the Weyl-squared term
@@
This result for the central charge $a$ is well known, as noted by Imbimbo et al. [43]. For the particular case, but comprising a large family of gravitational theories, of Einstein-like gravities (cf. [44]), there is further structure in the coefficient $\gamma$
@@
and, after basic algebraic steps, we obtain an expression for $c$ that effectively ’tweaks’ the standard Einstein-Hilbert result
$c\,=\,-h^{\prime}(f_{\infty})\cdot\dfrac{d\pi^{d/2}\cdot\tilde{L}^{d-1}}{[(d/2)!]^{2}\cdot l_{P}^{d-1}}\,=\,-h^{\prime}(f_{\infty})\cdot c^{E}|_{\tilde{L}}$. (56)
@@
This can be readily verified by taking the (log-)derivative of the holographic central charge $a$ (eqn.53)
@@
where we have used the on-shell condition to trade ${\cal L}(f_{\infty})$ by $\frac{2\,f_{\infty}}{d+1}\,{\cal L}^{\prime}(f_{\infty})$ and the fact that the latter only depends on $\tilde{L}$, as opposed to ${\cal L}(f_{\infty})$ that depends as well on $L$. That makes it possible to act with derivatives with respect to $f_{\infty}$. Using our universal relation, we end up with a prescription to compute $C_{T}$ directly from the type-A central charge $a$ in Einstein-like gravities. This ought to be compared with eq.(18) and the footnote [90] in [44].

### v2

# Universal relation between $C_{T}$ and the CFT Weyl anomaly

@@ Abstract
We establish a universal relation between the coefficient $C_{T}$ of the energy–momentum tensor two‑point function and the coefficient $c$ multiplying the term quadratic in the Weyl tensor in the Weyl anomaly of a generic even‑dimensional conformal field theory. Our first derivation combines long‑known holographic results for $C_{T}$ and for the Weyl anomaly in Einstein bulk gravity with a recently obtained Chern–Gauss–Bonnet formula for compact Einstein manifolds. This theorem isolates the Weyl‑squared contribution in the relation between the Euler density and the $Q$-curvature, allowing us to identify the relevant quadratic term unambiguously. We then provide two genuine CFT derivations based on the renormalization‑group running of the TT correlator with respect to the arbitrary but necessary mass scale $\mu$. Several known examples are revisited to illustrate and validate the general result.
@@ I Introduction
In even spacetime dimensions, CFTs exhibit a trace (or Weyl or conformal) anomaly when coupled to a background metric [2]. Although the classical theory is Weyl invariant, the quantum effective action acquires a non‑vanishing trace proportional to local curvature invariants. The anomaly takes the schematic form [3, 4, 5]
$(4\pi)^{d/2}\langle T\rangle=-a\,E_{d}+\sum_{i}c_{i}I_{i}+ttd,$ (2)
where the Euler density $E_{d}$ carries the type-A central charge $a$, the pointwise Weyl invariants $I_{i}$ carry the type-B central charges $c_{i}$, and ttdttd stands for trivial anomalies or trivial total derivatives whose conformal primitives are local curvature invariants. The central charges, traditionally denoted $a$ and $c$ in four dimensions and by suitable generalizations in higher even dimensions, are intrinsic CFT data.
@@
  • d=2: There is only one central charge, $c$, and it is defined by
    $4\pi\langle T\rangle=\frac{c}{6}R.$
@@
    $\langle T\rangle=-a\,E_{6}+c_{1}I_{1}+c_{2}I_{2}+c_{3}I_{3}+ttd,$
    where $I_{1}$ and $I_{2}$ are cubic in the Weyl tensor, while $I_{3}$ contains terms of the schematic form $W\nabla^{2}W+lot+ttd$. The lower-order terms in derivatives (lotlot) clearly do not contribute to the quadratic expansion around flat space and are not directly connected to $C_{T}$. Therefore, it is the central charge $c_{3}$ that relates to $C_{T}$ [1, 7]
@@
    In holographic Einstein gravity, where the bulk action is just the Hilbert-Einstein action plus a negative cosmological constant, the coefficients of the holographic Weyl anomaly [8, 9] are all determined by a single bulk coupling: that of the critical $Q$-curvature [10]. That enforces linear relations among the various $c_{i}$ and $a$, but structurally, $C_{T}$ still tracks the unique quadratic-in-$W$ direction in the space of invariants, just as in 6d it tracks $c_{3}$. In $8d$, Chen and Lü [11] (see also [14, 13]) have recently found a quadratic in Weyl invariant $I^{(8)}_{10}$ whose coefficient $c_{10}$ becomes equal to $a$ in holographic Einstein gravity. Comparing with the long-known result of Liu and Tseytlin for $C_{T}$ [12], they obtain (in our conventions)
@@
    where we have chosen a particular pointwise Weyl invariant $W^{(d)}_{2,1}$, recently introduced by Case et at. in [13] that descends from the ambient $\left(\widetilde{\nabla}^{2}\right)^{d/2-2}\widetilde{W}^{2}$ in the Fefferman–Graham program. Restricted to Einstein manifolds, it enters an explicit expression for the Chern–Gauss–Bonnet formula that allows for the interchange between the Euler density and the $Q$-curvature. With this insight, we tune the coefficient $c$ in the anomaly to enforce the equality $c=a$ when the anomaly becomes just the critical $Q$-curvature.
    Alternatively, we are also free to choose instead a different Weyl invariant quadratic in the Weyl tensor that naturally extends the $W^{2}$ and the $I_{3}$ from 4d and 6d, respectively (see e.g. [15]). The numerical coefficient is obtained by the expansion of the $Q$-curvature to terms quadratic in curvature (see subsection II.C below) to enforce the equality $c=a$ for the $Q$-curvature,
@@
In what follows we will explain how to obtain the following universal relation between $C_{T}$ and $c$, that accommodates the low dimensions and extrapolates to any even higher dimension [^1] Dimension two is special, our general result can be modified to include this particular case as well by absorbing a factor six in the definition of the 2d $c$, setting instead $4\pi\langle T\rangle=c\,R$
$C_{T}=\dfrac{d}{d-1}\cdot\dfrac{(d+1)!}{(d/1)!}\cdot c\qquad\text{ for even }d>2.$ (5)
Alternatively, in terms of the critical $Q$-curvature $Q_{d}$ and the coefficient $C_{TT}$ of the TT-correlator in momentum space  [^2] We thank P. McFadden for pointing out the suitability of the $C_{TT}$ coefficient, as opposed to the more conventional $C_{T}$., the explicit dimension dependence can be absorbed and the universality appears in a surprisingly simple form
$\boxed{\langle T\rangle\,=\,2\,C_{TT}\cdot Q_{d}+lot+ttd,}$ (6)
where the lower-order terms in derivatives (lotlot) now include the Euler density and only the pointwise Weyl invariants of the type-B trace anomaly, which are cubic and higher in curvature.
Understanding these relations is essential for characterizing the space of consistent CFTs in higher dimensions and to connect flat‑space observables with geometric responses. The present work develops this connection in full generality for even‐dimensional CFTs. We follow a holographic approach and combine the long-known holographic results for $C_{T}$ [12] and Weyl anomaly [8],[9],[17] for bulk Einstein gravity, together with a recently derived Chern-Gauss-Bonnet theorem on a compact Einstein manifold that allows one to pinpoint the term quadratic in the Weyl tensor in the relation between Euler density and $Q$-curvature. We also confirm our findings through two purely CFT computations that exploit the fact that the trace anomaly determines the scaling behavior of the CFT partition function and effective action.
We start in Section II.1, by reviewing the holographic computation of $C_{T}$ and relating it to the bulk Newton’s constant, or, equivalently, the Planck length. In II.2, we discuss the holographic Weyl anomaly for Einstein gravity with a negative cosmological constant and its relation with the $Q$-curvature. In II.3, we present the Chern-Gauss-Bonnet theorem on compact even-dimensional Einstein manifolds, as recently obtained by Case et al. [13], which relates the Euler density to the $Q$-curvature and pointwise Weyl invariants. In particular, we keep track of the coefficient $c$ of the quadratic term in the Weyl tensor. In II.4, we combine the previous ingredients to obtain the universal relation between $C_{T}$ and $c$. In Section III, we present two alternative derivations within the CFT. Finally, in Section IV, we compare the universal relation with several known examples, illustrate them, and conclude in Section V.
@@ II.3 Weyl-squared content of the $Q$-curvature
$\mbox{Pf}=(2\lambda)^{d/2}\cdot(d-1)!!\,+\,\dfrac{(-2)^{2-d/2}}{(d/2-1)!}\cdot\frac{1}{8}\,W^{(d)}_{2,1}+lot+ttd.$ (13)
@@
$E_{d}=\dfrac{(-)^{d/2}}{2^{d/2}}\,\epsilon_{d}\,\epsilon_{d}\,Riem...Riem\qquad\text{ and }\qquad\mbox{Pf}=\dfrac{(d-1)!!}{2^{d/2}\,d!}\,\epsilon_{d}\,\epsilon_{d}\,Riem...Riem\,=\,\dfrac{(-)^{d/2}\,(d-1)!!}{d!}\cdot E_{d}~.$ (14)
@@
We are now in a position to write down the holographic value for $c$ from eqn.12
@@
$Q_{d}=-\,\dfrac{d-2}{8(d-3)}\,W(-\nabla^{2})^{d/2-2}W+lot+ttd.$ (18)
@@
When written in terms of the $C_{TT}$ coefficient of the TT-correlator in momentum space ([55])
$\langle\langle T_{ab}({\bf p})\,T_{cd}(-{\bf p})\rangle\rangle\,=\,C_{TT}\,\Pi_{ab,cd}({\bf p})\,p^{d},$ (20)
with $\Pi_{ab,cd}$ being the transverse-traceless projector and the Dirac delta enforcing momentum conservation factored out, the above relation simplifies to
$\boxed{C_{TT}=-\frac{(-1)^{d/2}\,d}{2\,(4\pi)^{d/2}}\cdot c}~.$ (21)
@@
III Universal relation between $C_{T}$ and $c$ : CFT derivations
The CFT derivations exploit the fact that the trace anomaly controls the scale variation of the partition function $Z_{CFT}$ or effective action $W_{CFT}$ (see e.g. [1])
@@ III Universal relation between $C_{T}$ and $c$ : CFT derivations
On the RHS, in turn, we need the kernel corresponding to the second metric variation of the trace anomaly. It suffices to keep track of the quadratic term in the Weyl tensor. Furthermore, the Euler term does not contribute. One can directly expand the term quadratic in Weyl up to second order in the metric perturbation around flat space, or trade it for the Euler term and $Q$-curvature modulo total derivatives. The $Q$-curvature is pure Ricci and can be easily varied in 4D and 6D, where explicit expressions for it are at hand. Cubic and higher-order terms in the Weyl tensor can safely be discarded when expanding around flat space in the computation of the two-point function, since their contribution to the effective action is cubic or higher in the metric perturbation.
@@
that must be cancelled by the one-loop counterterm, which in turn, is given by the trace anomaly (see e.g. [25], eqns. 21-26 therein for 2D and 4D  [^6] We are grateful to M.J. Duff for information on this point. This same pole matching led to the discovery of the GJMS operators in the two-point function for scalar operators in [53]. See also  [54] for a thorough analysis of scalar operators.).
@@ III.0.4 Alternative derivation in 6D
After some algebraic manipulations while expanding the Ricci tensor and scalar, as expected, one obtains the very same relation as before  [^7] This 6D relation was announced by Sen and Sinha [27], but it contained a wrong numerical coefficient in the holographic computation of the central charge $c_{3}$ that spoiled the relation with $C_{T}$ in the journal version. The arXiv version 4, in turn, contains the correct result that agrees with the holographic expectation; see eqns. 3.22 and D.29 therein. We are grateful to Felipe Diaz for clarification of this issue..
@@ III.0.6 General even $D$: $Q$-curvature and GJMS-like operators
But there is a remarkable observation following this line of reasoning: the second metric variation of the $Q$-curvature, which can be thought of as a particular conformal gravity Lagrangian admitting Einstein solutions, is just the kinetic term of the corresponding Weyl graviton when restricted to the transverse-traceless component of the metric fluctuation. This is nothing but the GJMS-like operator on symmetric transverse-traceless two-tensors that happens to factorize into shifted Lichnerowicz Laplacians on Einstein manifolds, as shown some time ago by Matsumoto [33]. In particular, in flat space, it reduces to a power of the Laplacian. It is just a matter of keeping track of the overall coefficient to be able to obtain the numerical factor that fixes the RHS of eqn.32999We are grateful to A.A.Tseytlin for discussion of this point.
@@ IV Examples
11 $\partial^{2}$ conformal scalar $8/7$ 66
22 $\partial^{4}$ conformal scalar 32/7-32/7 24-24
33 $\partial^{6}$ conformal scalar 2424 126126
44 $\partial^{8}$ conformal scalar 256-256 1344-1344
@@
$C^{E}_{T}=4\pi^{d/2}\cdot\frac{\Gamma(d+2)}{(d-1)\,\Gamma^{3}(d/2)}\cdot\frac{L^{d-1}}{l_{\text{P}}^{d-1}}\qquad vs.\qquad c^{E}=\dfrac{4\pi^{d/2}}{d\,\Gamma^{2}(d/2)}\cdot\dfrac{L^{d-1}}{l_{\text{P}}^{d-1}}~.$ (45)
@@
Let us follow a shortcut route [42, 43, 44] to get this algebraic equation . This is equivalent to extremizing the gravitational action with respect to the putative AdS radius as in [36]. We exploit the scale invariance of the action to write
@@
$S_{\text{G}}[\mu^{2}\hat{g}_{AdS}]=\mu^{d+1}\int_{\mathcal{M}}d^{d+1}x\sqrt{\hat{g}_{AdS}}\;{\cal L}\left(\frac{f_{\infty}}{\mu^{2}}\right),$ (49)
@@
We can now apply our simple holographic recipe to determine the holographic central charge $a$ by evaluating it on a particular Poincaré-Einstein bulk metric with an Einstein metric at its conformal infinity [44, 40]. From the pure-Ricci part, we get the action at $AdS$ precisely, and from the deviations, involving the Weyl tensor, it is enough to keep the Weyl-squared term
@@
This result for the central charge $a$ is well known, first obtained by Imbimbo et al. [46]. As for the $c$ charge, in the particular case of Einstein‑like gravities—though this still comprises a large family of gravitational theories (cf. [47])—the structure of the coefficient $\gamma$ admits further refinement
@@
and, after basic algebraic steps, we derive an expression for $c$ that tweaks the canonical Einstein–Hilbert value in a notably simple way
$c\,=\,-h^{\prime}(f_{\infty})\cdot\dfrac{d\pi^{d/2}\cdot\tilde{L}^{d-1}}{[(d/2)!]^{2}\cdot l_{\text{P}}^{d-1}}\,=\,-h^{\prime}(f_{\infty})\cdot c^{E}|_{\tilde{L}}$. (59)
@@
This can be readily verified by taking the (log-)derivative of the holographic central charge $a$ (eqn.56)
@@
where we have used the on-shell condition to trade ${\cal L}(f_{\infty})$ by $\frac{2\,f_{\infty}}{d+1}\,{\cal L}^{\prime}(f_{\infty})$ and the fact that the latter only depends on $\tilde{L}$, as opposed to ${\cal L}(f_{\infty})$ which also depends on $L$. That makes it possible to act with derivatives with respect to $f_{\infty}$. Using the present universal relation, we end up with a prescription to compute $C_{T}$ directly from the type-A central charge $a$ in Einstein-like gravities. A comparison with [47] (eqn.18 and footnote [90] therein) is also in order.